The 1st term of an arithmetic sequence is 6 times the 5th term. The sum of the 2nd and 8th terms is 8. Find the 20th term of this sequence?
Answer by richwmiller(17219)

a =6(a + 4d)
a =6a + 24d
a2 = a + d
a8 = a + 7d
8=2a+8d
4=a+4d
24=6a+24d
a =6a + 24d
a=24
24=144+24d
-120=24d
-5=d
an = a + (n - 1)d
a20 = 24 -5(19)
a20=-71